A soap, C17H33CO2Na, has a molecular weight of 304 g/mole. 1 kg of this soap is discharged into a pond containing 1.8 x 10° L. 5. Write the reaction between this soap and O2 Assuming that the soap is completely degraded to CO2 and water within 5 days, compute the BOD. Express your answer in mg O2/L. a. b.
Answer #1

a) C17H33CO2Na + 97/2O2>>>>> 17CO2 + 33/2H2O +Na-

b) So for BOD , required amount of O2 = 97/2*32 = 1552gm

BOD = 1552*103/1.8*106 (mg/L) = 8.62*10-1(mg/L)